# An Idiosyncrasy of Time-discretization in Reinforcement Learning

### v1

@@ 1 Introduction
Reinforcement learning provides a framework for solving sequential decision making problems based on evaluative feedback (Sutton & Barto, 2018). It remains a promising approach for robot learning as it can allow for real-time adaptation of behavior. Many reinforcement learning algorithms assume that the agent-environment interaction occurs at synchronous, discrete time steps, where the environment waits for an action before advancing. In contrast, real-world physical systems are continuous in time, and do not wait for an agent’s input. As such, time-discretization becomes a necessary and important consideration (Mahmood et al., 2018a).
Prior work suggests that current reinforcement learning algorithms are sensitive to the choice of discretization. Tallec et al. (2019) emphasizes that action-values converge to state-values as the discretization interval approaches zero, creating degenerate cases for algorithms like Q-learning. Similarly, Munos (2006) shows that the variance of policy gradients can explode under the same limit. Mahmood et al. (2018a) details the trade-off between having fine-grained control and being able to discern the changes between subsequent states. Finally, Farrahi & Mahmood (2023) provides guidelines for time-discretization-aware parameter selection by acknowledging how changes in discrete-time parameters influence the underlying continuous-time objective.
In this work, we explicitly view the discrete-time objective as a discrete approximation of the continuous-time objective. By considering when rewards occur, particularly in existing continuous-control environment setups, we identify an idiosyncratic dependence on the choice of discretization beyond those listed in Tallec et al. (2019) and Farrahi & Mahmood (2023). Specifically, the discrete-time return can be viewed as a mixture of two Riemann sums. We characterize and demonstrate that this is a relatively poor integral approximation in comparison with a conventional Riemann sum, and provide a simple modification to the definition of the return to better align the objectives.
@@
  • Acknowledgement and characterization of a discrepancy when naively applying a discrete-time reinforcement learning algorithm to a discretized continuous-time environment.
  • A simple modification to the definition of the return to avoid a nuanced dependence on time-discretization, based on an integral approximation perspective.
  • Characterization of when the modification is most impactful, supported by empirical results demonstrating improved alignment with an underlying continuous time objective.
@@ 2 Definitions of the Return
Reinforcement learning (Sutton & Barto, 2018) is a framework for sequential decision making from evaluative feedback. A learning system, denoted an agent, observes its current situation and selects an action, after which it observes a new situation while receiving a reward. The agent’s objective is to learn to act so as to maximize its expected return- a discounted sum of future rewards. In discrete-time (Sutton & Barto, 2018), the return from a time step $t$ onward is defined to be:
$\ddot{G}_{t}=\sum_{k=t}^{T-1}{\gamma^{k-t}R_{k+1}}$ (1)
where $T$ is the final time step in an episodic task, or $\infty$ in a continuing, infinite-horizon one. In continuous-time reinforcement learning (Doya, 2000; Mehta & Meyn, 2009; Frémaux et al., 2013; Lee & Sutton, 2021; Tallec et al., 2019), we instead define the integral return from time step $t$ onward:
@@
Such a formulation is pertinent to applications involving real-time interaction, such as robotics. Despite being continuous in time, they are often digitally controlled, and as such time-discretization becomes a necessary consideration (Mahmood et al., 2018a).
@@ 3 When Rewards Occur
There are notation inconsistencies in the literature with respect to time indices in the discrete time return (Equation 1). Some define it to start from $R_{t+1}$ (Sutton, 1988; Precup et al., 2000; van Seijen et al., 2009; Barreto et al., 2017), as presented in this document, while some would start from $R_{t}$ (Watkins, 1989; van Hasselt, 2010; Mnih et al., 2015; Wang et al., 2016). This inconsistency is inconsequential when solely considering the discrete-time setting, as the rewards occur at the same locations in an agent’s stream of experience. However, it has implications when viewed as a discrete approximation to an underlying integral return. Thus, it is worth considering when rewards occur.
We emphasize the focus on a setting where there is an underlying continuous-time objective, of which a digital learning agent samples at an arbitrary (and potentially variable) frequency. Despite the discrete-time notation inconsistencies, it is often agreed upon that from the agent’s perspective, the reward and next state are jointly observed. This is reflected in environment step calls in relatively standard reinforcement learning APIs (Brockman et al., 2016), agent-environment interaction diagrams (Sutton & Barto, 2018), or explicit acknowledgement that reward can be a function of state, action, and next state (Puterman, 1994). In real-time settings which do not wait for an agent’s input, actions typically take time to execute and have an influence, and so meaningful evaluative feedback must come after time $t$. Hardware limitations on sampling rates further impose an inherent delay in when a system can receive feedback for an action. In many existing robotics environments, where the considered setting is especially pertinent, rewards are often explicitly computed based on the next time step’s state information (Todorov et al., 2012; Brockman et al., 2016; Mahmood et al., 2018b), for example, rewards based on distance traveled in some direction between two time steps, or distance between an end-effector and a desired setpoint at the subsequent time step.
Of note, semi-MDPs and options (Sutton et al., 1999; Precup, 2000) address the problem of when rewards occur, but under the assumption that one has access to higher-frequency interaction with the environment to integrate the discounted sum of rewards within the discretization interval. It is akin to the agent being aware of and able to time when each component of a temporally-extended reward occurs. In this work, we consider when one does not have access to these higher-frequency samples, but is aware of how much time has elapsed between discrete decision points. Acquiring such information may not be possible due to hardware limitations, and highlights a nuance that arises when naively applying a discrete-time algorithm to a discretized continuous-time environment.
@@ 4 Implications for Time Discretization
If we consider rewards jointly arriving with the next state, at least from the agent’s perspective, this results in an idiosyncrasy with respect to approximating an underlying integral return. While the discrete-time returns may use inconsistent reward time-indices, they are consistent on when discounting begins: the first reward is given weight $\gamma^{0}=1$, with subsequent rewards weighted by increasing powers of $\gamma$. We can view the integral return in Equation 2 to be of the form:
@@
where $f(\tau)$ is the discounting term, and $g(\tau)$ is the reward signal. A right-point Riemann sum approximation to this would yield:
$\sum_{i=0}^{n-1}{f(\tau_{i})g(\tau_{i})\Delta}$ (4)
where $\Delta=\frac{T-t}{n}$ and $\tau=\{t+\Delta,t+2\Delta,...,T\}$. The right-point Riemann sum beginning with $t+\Delta$ aligns with an agent jointly receiving a reward with the observation of the next state. However, this sum would weight the first reward by $\gamma^{\Delta}\neq\gamma^{0}$. This highlights that if one naively applies a discrete-time reinforcement learning algorithm to a discretized continuous-time environment, it is akin to a left-point Riemann sum for discounting, and a right-point Riemann sum for rewards:
$\sum_{i=0}^{n-1}{f(\tau_{i})g(\tau_{i+1})\Delta}$ (5)
where $\tau\in\{t,t+\Delta,t+2\Delta,...,T\}$. See Figure 1 for a visualization of this Riemann sum. This sum still converges to the correct integral as $n\rightarrow\infty$, as Bliss’s Theorem (1914) allows each function to be evaluated at any point within the interval. However, for the specific case where a left-point Riemann sum is used for discounting, we expect this to perform worse than committing to a right-point Riemann sum. Due to the curvature of exponential decay, if one drew a rectangle with opposite corners at any two points, there will always be more area above the curve than below, implying an underestimate has strictly lower error than an overestimate. This is visualized in Figure 2.
@@
To rectify this discrepancy and commit to a right-point Riemann sum approximation, one would simply multiply the discrete-time return by a factor of $\gamma$ (assuming $\Delta=1$):
$\gamma\ddot{G}_{t}=\gamma R_{t+1}+\gamma^{2}R_{t+2}+\cdots$ (6)
For a fixed, pre-specified action cycle-time $\Delta$, there is no loss of generality as the discrete-time return is proportional by a factor of $\gamma^{\Delta}\Delta$. However, this is not the case when $\Delta$ may vary over time, e.g., due to an adaptive algorithm (Karimi et al., 2023) or inherent stochasticity. These concerns similarly apply to a variable $\gamma$ and may extend toward tuning fixed-$\Delta$ and/or $\gamma$ in practice in terms of a nuanced and unintuitive dependence on discretization. See appendix A for a discrete-time return visualization with variable interval sizes. To emphasize the dependence on $\Delta$, we note the explicit right-point Riemann sum return:
@@
$=\gamma^{\Delta_{t+1}}R_{t+1}\Delta_{t+1}+\gamma^{\Delta_{t+1}+\Delta_{t+2}}R_{t+2}\Delta_{t+2}+\cdots$
Prior work has acknowledged the modifications of (1) scaling rewards by $\Delta$, and (2) exponentiating $\gamma$ by $\Delta$ (Tallec et al., 2019; Farrahi & Mahmood, 2023) in improving robustness to time-discretization. The key difference and contribution in Equation 7 being the earlier discounting.
@@ 5 Comparison with Standard Riemann Sums
To see how the discrete-time return (DTR) in Equation 5 fares against a right-point Riemann sum, we evaluate them on randomly generated continuous-time signals. Inspired by robotics, we consider periodic signals, and Gaussian mixtures. Periodic signals are comparable to signals pertaining to robot locomotion, while Gaussian mixtures instead resemble both sparse and distance-based rewards, depending on the spread of each Gaussian. We fix the signal length to 3 seconds, with no loss of generality due to being continuous in time. Each signal generator is detailed below:
@@
Random Gaussian Mixtures - This signal sums 6 Gaussians $\sum_{i=0}^{5}{\mathcal{N}(\mu_{i},\sigma_{i})}$ with means $\mu_{i}\sim\mathcal{U}(0,3)$, and standard deviations $\sigma_{i}\sim\mathcal{U}(0,\frac{3}{2})$.
For each method, we varied the number of intervals $n\in\{5,10,25,50,100\}$, the discount factor $\gamma\in\{0.5,0.75,0.875\}$, and measured the absolute error of the integral approximation- compared against a mid-point Riemann sum with $10^{4}$ intervals. The values of $\gamma$ used may appear small and unrepresentative of typical values. We however note that the discount is per second, and that for a robot sampling every 30 ms, $\gamma=0.5$ is effectively $\gamma^{\Delta}=0.5^{0.03}\approx 0.98$ per discrete time step. Averaged across $10^{6}$ randomly generated signals of each type, the results can be seen in Figure 3.
@@
As expected, the errors generally increase as $\Delta\propto\frac{1}{n}$ increases. There’s a consistent dip in error with the periodic signals, likely due to the intervals coincidentally aligning with the pre-specified frequencies. Across all settings, DTR had larger absolute error, consistent with our hypothesis that DTR would perform worse than right-point on discounted signals. The gap closes as $\gamma\rightarrow 1$, as the sums are equivalent at this extreme.
We then considered stochastic intervals to simulate variable time-discretization. This was implemented by sampling, sorting, and re-scaling a set of $n+1$ uniform random points representing interval endpoints. This is particularly pertinent as DTR is no longer proportional to right-point, and reflects the variability in applications on real-time systems. Fixing $\gamma=0.75$, Figure 4 shows results averaged across $10^{6}$ randomly generated signals of each type, plotted against average $\Delta$. Errors generally increased, with DTR maintaining larger approximation error across every setting.
@@
Lastly, to see whether results hold beyond exponential discounting, we considered the product of each pair of the signal generators. This evaluates each sum in a more general numerical integration setting, while having implications for variable, transition-dependent $\gamma$ in reinforcement learning. Averaged across $10^{6}$ randomly generated signal pairs, the results can be seen in Figure 5. Perhaps surprisingly, the gap between DTR and the right-point Riemann sum widens dramatically. These results suggest that beyond the structure of discounting, DTR is a generally worse integral approximation.
@@
A key takeaway from these results is that shifting the discount factor in the discrete-time return yields a better prediction target (e.g., in value-based methods) in terms of error between the integral return. To reiterate, in the fixed $\Delta$ case, the sums are proportional despite the gaps in approximation error. This suggests that the improvement is inconsequential for control. However, such improved alignment is expected to improve control performance in the variable $\Delta$ setting, in terms of maximizing the underlying integral return. We explore this further in the next section.
@@ 6 Discretized Continuous-time Control
To evaluate the right-point Riemann sum in a continuous-time control setting, we build off of the REINFORCE (Williams, 1992) algorithm. Such a choice is due to the algorithm’s simplicity, allowing for more confidence in attributing differences in performance. We specifically use online REINFORCE with eligiblity traces (Kimura et al., 1995) and dropped $\gamma^{t}$ term, summarized by:
@@
$\mathbf{z}$ $\leftarrow\gamma^{\Delta_{t+1}}\mathbf{z}$
where $\Delta_{t+1}$ is the elapsed time between time steps $t$ and $t+1$, $R_{eff}=R_{t+1}\Delta_{t+1}$ for the discrete-time return, and $R_{eff}=\gamma^{\Delta_{t+1}}R_{t+1}\Delta_{t+1}$ for the right-point Riemann sum. The above algorithm employs the recommendations of (Farrahi & Mahmood, 2023) for making algorithms more robust to time-discretization, emphasizing that the proposed right-point modification is complimentary. Each agent’s policy used a two-hidden-layer fully-connected network with $tanh$ activations, with its output being treated as the mean of a Gaussian with an initial (bias unit) standard deviation of 1.
@@
The environment fixed $\gamma=0.25$, which when using an interval of 40 ms, corresponds with discrete-time $\gamma^{0.04}\approx 0.95$. We considered target discretization intervals $\Delta_{\mu}\in\{40,80,120\}$ ms with a 4 second time limit, and measured the episodic integral return. Averaged over 100 25-minute runs, Figure 6 shows parameter sensitivity curves, as well as learning curves under the best parameters.
@@
An initial observation is a systematic “lag” between the two algorithms in the sesntivity curves at low $\alpha$. This is due the return magnitudes being roughly proportional by a factor of $\mathbb{E}[\gamma^{\Delta_{t}}]$. If one absorbs this factor into the step-size, the right-point Riemann sum can be viewed as using a smaller effective $\alpha$ in the policy gradient update. Scaling the figure to use this effective $\alpha$ can be found to well-align the curves at low $\alpha$. Nevertheless, we find that after accounting for this shift, REINFORCE with the right-point Riemann sum performs better for large $\alpha$, and can significantly outperform the discrete-time return with both algorithms properly tuned. The right-point Riemann sum is seen to improve with increasing $\Delta_{\mu}$, in line with the approximation error results in Section 5. Acknowledging that the two returns are roughly proportional by $\mathbb{E}[\gamma^{\Delta_{t}}]$, improvements are expected as this term deviates from 1, i.e., decreasing $\gamma$ or increasing $\Delta_{\mu}$.

### v2

@@ 1 Introduction
Reinforcement learning provides a framework for solving sequential decision making problems based on evaluative feedback (Sutton and Barto,, 2018). It remains a promising approach for robot learning as it can allow for real-time adaptation of behavior. Many reinforcement learning algorithms assume that the agent-environment interaction occurs at synchronous, discrete time steps, where the environment waits for an action before advancing. In contrast, real-world physical systems are continuous in time, and do not wait for an agent’s input. As such, time-discretization becomes a necessary and important consideration, as evidenced by Mahmood et al., 2018a .
Prior work suggests that current reinforcement learning algorithms are sensitive to the choice of discretization. Tallec et al., (2019) emphasize that action-values converge to state-values as the discretization interval approaches zero, creating degenerate cases for algorithms like Q-learning. Similarly, Munos, (2006) showed that the variance of policy gradients can be infinite under the same limit. Zhang et al., (2023) characterize a fundamental bias-variance trade-off with the degree of discretization while Mahmood et al., 2018a detail another trade-off between having fine-grained control and being able to discern the changes between subsequent states. Finally, Farrahi and Mahmood, (2023) provide guidelines for time-discretization-aware parameter selection by acknowledging how changes in discrete-time parameters influence the underlying continuous-time objective.
In this work, we explicitly view the discrete-time objective as a discrete approximation of the continuous-time objective. By considering when rewards occur, particularly in existing continuous-control environment setups, we identify an idiosyncratic dependence on the choice of discretization beyond those listed by Tallec et al., (2019) and Farrahi and Mahmood, (2023). Specifically, the discrete-time return can be viewed as mixing two Riemann sums. We characterize and demonstrate that this is a relatively poor integral approximation in comparison with a conventional Riemann sum and provide a simple modification to the definition of the return to better align the objectives.
@@
  • We acknowledge and characterize an issue with naively applying a discrete-time reinforcement learning algorithm to a discretized continuous-time environment in terms of a discrepancy between the discrete-time and continuous-time definitions of the return.
  • Based on an integral approximation perspective, we propose a simple modification to the definition of the return to alleviate this idiosyncratic dependence on time-discretization.
  • We characterize when the modification will have a modest impact and support our claim with empirical evaluation in both continuous-time prediction and control.
@@ 2 Definitions of the Return
In discrete-time reinforcement learning, the discounted return from time step $t$ onward is defined as:
$G_{t}=\sum_{k=t}^{T-1}{\gamma^{k-t}R_{k+1}},$ (1)
where $T$ is the final time step of an episodic task, or $\infty$ in an infinite-horizon setting. In continuous-time reinforcement learning (e.g., Doya,, 2000; Mehta and Meyn,, 2009; Frémaux et al.,, 2013; Lee and Sutton,, 2021; Tallec et al.,, 2019), we instead define the integral return from time step $t$ onward:
@@
This formulation is pertinent to applications with real-time interaction (e.g., robotics). Despite being continuous in time, robots are often digitally controlled, necessitating understanding the impact of the choice of time-discretization and how it relates these two objectives.
@@ 3 When Rewards Occur
There are notational differences in the literature with respect to time indices in the discrete time return (Equation 1). Some define it to start from $R_{t+1}$ (e.g., Sutton,, 1988; Precup et al.,, 2000; van Seijen et al.,, 2009; Barreto et al.,, 2017), as presented in this document, while some would start from $R_{t}$ (e.g., Watkins,, 1989; van Hasselt,, 2010; Mnih et al.,, 2015; Wang et al.,, 2016). This inconsistency is inconsequential when solely considering the discrete-time setting as the rewards occur at the same locations in an agent’s stream of experience. However, it has implications when viewed as a discrete approximation to an underlying integral return. Thus, it is worth considering when rewards occur.
We emphasize the focus on a setting where there is an underlying continuous-time objective of which a digital learning agent samples at an arbitrary (and potentially variable) frequency. Despite the discrete-time notational differences, it is often agreed upon that from the agent’s perspective, the reward and next state are jointly observed. This is reflected in environment step calls in relatively standard reinforcement learning APIs (e.g., Brockman et al.,, 2016), agent-environment interaction diagrams (e.g., Sutton and Barto,, 2018), or explicit acknowledgement that reward can be a function of state, action, and next state (e.g., Puterman,, 1994). In real-time settings that do not wait for an agent’s input, meaningful evaluative feedback must come after time $t$ as actions take time to execute and to have a causal influence. Hardware limitations on sampling rates further delay when a system can receive feedback for an action. In many existing robotics environments, where the considered setting is especially pertinent, rewards are often explicitly computed based on the next time step’s state information. For example, rewards based on distance traveled in some direction between two time steps, or distance between an end-effector and a desired setpoint at the subsequent time step, as done by Todorov et al., (2012), Brockman et al., (2016), and Mahmood et al., 2018b .
Of note, semi-MDPs and options (Sutton et al.,, 1999; Precup,, 2000) address the problem of when rewards occur, but under the assumption that one has access to higher-frequency interaction with the environment to integrate the discounted sum of rewards within the discretization interval. It is akin to the agent being aware of and able to time when each component of a temporally-extended reward occurs. Here, we consider when one does not have access to these higher-frequency samples but is aware of how much time has elapsed between discrete decision points. Acquiring such information may not be possible due to hardware limitations, and highlights a nuance that arises when naively applying a discrete-time algorithm to a discretized continuous-time environment.
@@ 4 Implications for Time Discretization
If we consider rewards jointly arriving with the next state, at least from the agent’s perspective, then there is an idiosyncrasy with respect to approximating an underlying integral return. While definitions of the discrete-time return may differ in their use of reward time indices, they are consistent on when discounting begins: the first reward is given weight $\gamma^{0}=1$, with subsequent rewards weighted by increasing powers of $\gamma$. We can view the integral return in Equation 2 to be of the form:
@@
where $f(\tau)$ is the discounting term and $g(\tau)$ is the reward signal. A right-point Riemann sum approximation to this would yield:
$\sum_{i=0}^{n-1}{f(\tau_{i})g(\tau_{i})\Delta},$ (4)
where $\Delta=\frac{T-t}{n}$ and $\tau=\{t+\Delta,t+2\Delta,...,T\}$. The right-point Riemann sum beginning with $t+\Delta$ aligns with an agent jointly receiving a reward with the observation of the next state. However, this sum would weight the first reward by $\gamma^{\Delta}\neq\gamma^{0}$. This highlights that if one naively applies a discrete-time reinforcement learning algorithm to a discretized continuous-time environment, it is akin to a left-point Riemann sum for discounting and a right-point Riemann sum for rewards:
$\sum_{i=0}^{n-1}{f(\tau_{i})g(\tau_{i+1})\Delta},$ (5)
where $\tau\in\{t,t+\Delta,t+2\Delta,...,T\}$. See Figure 1 for a visualization of this Riemann sum. This sum still converges to the correct integral as $n\rightarrow\infty$ as Bliss’s Theorem (1914) shows that each function may be evaluated anywhere in the interval. However, for the specific case where a left-point Riemann sum is used for discounting, we expect this to perform worse than committing to a right-point Riemann sum. If one draws a rectangle with opposite corners at any two points of an exponential decay, the area above and below the curve represents the approximation errors of left- and right-point Riemann sums, respectively. There will always be more area above the curve than below due to the curvature of exponential decay, implying that an underestimate (right-point) has strictly lower error than an overestimate (left-point). This is visualized in Figure 2.
@@
To rectify this discrepancy and commit to a right-point Riemann sum approximation, we simply multiply the discrete-time return by a factor of $\gamma$. For example, with $\Delta=1$:
$\gamma G_{t}=\gamma R_{t+1}+\gamma^{2}R_{t+2}+\cdots.$ (6)
For a fixed, pre-specified action cycle-time $\Delta$, there is no loss of generality, as the discrete-time return is proportional by a factor of $\gamma^{\Delta}\Delta$. However, this is not the case when $\Delta$ may vary over time, for example, due to an adaptive algorithm (e.g., Karimi et al.,, 2023) or inherent stochasticity. These concerns similarly apply to a variable $\gamma$ and may extend toward tuning fixed-$\Delta$ or $\gamma$ in practice in terms of an unintuitive dependence on discretization. To emphasize the dependence on $\Delta$, we note the more explicit definition of the right-point Riemann sum return:
@@
$=\gamma^{\Delta_{t+1}}R_{t+1}\Delta_{t+1}+\gamma^{\Delta_{t+1}+\Delta_{t+2}}R_{t+2}\Delta_{t+2}+\cdots.$
Tallec et al., (2019) and Farrahi and Mahmood, (2023) have acknowledged the modifications of scaling rewards by $\Delta$ and exponentiating $\gamma$ by $\Delta$ in terms of improving robustness to time-discretization. The key difference and contribution in Equation 7 being the earlier discounting.
@@ 5 Comparison with Standard Riemann Sums
To see how the discrete-time return (DTR) in Equation 5 compares with a right-point Riemann sum, we use them to numerically integrate random continuous-time signals. Inspired by robotics, we consider periodic signals and Gaussian mixtures. Periodic signals are comparable to signals pertaining to robot locomotion, while Gaussian mixtures instead resemble both sparse and distance-based rewards depending on the spread of each Gaussian. We fix the signal length to 3 seconds, with no loss of generality due to being continuous in time. Each signal generator is detailed below:
@@
Random Gaussian Mixtures - This signal sums 6 Gaussians $\sum_{i=0}^{5}{\mathcal{N}(\mu_{i},\sigma_{i})}$ with means $\mu_{i}\sim\mathcal{U}(0,3)$ and standard deviations $\sigma_{i}\sim\mathcal{U}(0,\frac{3}{2})$.
For each method, we varied the number of intervals $n\in\{5,10,25,50,100\}$, the discount factor $\gamma\in\{0.5,0.75,0.875\}$, and measured the absolute error of the integral approximation relative to a fine-grained mid-point Riemann sum with $10^{4}$ intervals. The values of $\gamma$ used may appear small and unrepresentative of typical values. We however note that the discount is per second and that for a robot sampling every 30 ms, $\gamma=0.5$ is effectively $\gamma^{\Delta}=0.5^{0.03}\approx 0.98$ per discrete time step. Averaged across $10^{6}$ randomly generated signals of each type, the results can be seen in Figure 3.
@@
As expected, the errors generally increase as $\Delta\propto\frac{1}{n}$ increases. There is a consistent dip in error with the periodic signals which is likely due to the intervals coincidentally aligning with the pre-specified frequencies. Across all settings, DTR had larger absolute error and is consistent with our hypothesis that DTR would perform worse than right-point when integrating discounted signals. The gap closes as $\gamma\rightarrow 1$ as the sums are equivalent at this extreme.
We then considered stochastic intervals to simulate variable time-discretization. This was implemented by sampling, sorting, and re-scaling a set of $n+1$ uniform random points to represent interval endpoints. This is particularly pertinent as DTR is no longer proportional to right-point and reflects the variability in applications on real-time systems. Fixing $\gamma=0.75$, Figure 4 shows results averaged across $10^{6}$ randomly generated signals of each type plotted against average $\Delta$. Errors generally increased, with DTR maintaining larger approximation error across every setting.
@@
Lastly, to see whether results hold beyond exponential discounting, we considered the product of each pair of the signal generators. This evaluates each sum in a more general numerical integration setting, while resembling transition-dependent $\gamma$ as White, (2016) has advocated for in reinforcement learning. Averaged across $10^{6}$ randomly generated signal pairs, the results can be seen in Figure 5. Perhaps surprisingly, the gap between DTR and the right-point Riemann sum widens dramatically. This suggests that beyond the structure of discounting, DTR is a generally worse integral approximation.
@@
A key takeaway from these results is that shifting the discount factor in the discrete-time return yields a better prediction target (e.g., in value-based methods) in terms of error between the integral return. To reiterate, in the fixed $\Delta$ case, the sums are proportional despite the gaps in approximation error. This suggests that the improvement is inconsequential for control. However, in the variable $\Delta$ setting, we expect that learning from estimates which better approximate the underlying integral return should improve the capability to maximize it. We explore this further in the next section.
@@ 6 Discretized Continuous-time Control
To evaluate the right-point Riemann sum in a continuous-time control setting, we build off of the REINFORCE algorithm (Williams,, 1992). Such a choice is due to the algorithm’s simplicity, allowing for more confidence in attributing differences in performance. We specifically use online REINFORCE with eligiblity traces (Kimura et al.,, 1995) and dropped the $\gamma^{t}$ term:
@@
$\mathbf{z}$ $\leftarrow\gamma^{\Delta_{t+1}}\mathbf{z},$
where $\Delta_{t+1}$ is the elapsed time between time steps $t$ and $t+1$, $R_{eff}=R_{t+1}\Delta_{t+1}$ for the discrete-time return, and $R_{eff}=\gamma^{\Delta_{t+1}}R_{t+1}\Delta_{t+1}$ for the right-point Riemann sum. The above algorithm employs the recommendations of Farrahi and Mahmood, (2023) for making algorithms more robust to time-discretization, emphasizing that the proposed right-point modification is complimentary.
@@
Each agent’s policy used a two-hidden-layer fully-connected network with $tanh$ activations, with its output being treated as the mean of a Gaussian with an initial (bias unit) standard deviation of 1. We fixed $\gamma=0.25$, which when using an interval of 40 ms, corresponds with $\gamma^{0.04}\approx 0.95$ per discrete time step. We considered target discretization intervals $\Delta_{\mu}\in\{40,80,120\}$ ms with a 4 second time limit and measured the episodic integral return. Averaged over 100 runs of 25 (simulation) minutes, Figure 6 shows parameter sensitivity curves and the best parameters’ learning curves.
@@
An initial observation is a systematic lag between the sensitivity curves of the two algorithms at low $\alpha$. This is due to the return magnitudes being roughly proportional by a factor of $\mathbb{E}[\gamma^{\Delta_{t}}]$. If one absorbs this factor into the step-size, the right-point Riemann sum can be viewed as using a smaller effective $\alpha$ in the policy gradient update. Scaling the figure to use this effective $\alpha$ can be found to align the curves at low $\alpha$. Nevertheless, we find that after accounting for this shift, REINFORCE with the right-point Riemann sum never performed worse and can significantly outperform the discrete-time return with both algorithms properly tuned. The right-point Riemann sum is seen to improve with increasing $\Delta_{\mu}$, in line with the approximation error results in Section 5. Acknowledging that the two returns are roughly proportional by $\mathbb{E}[\gamma^{\Delta_{t}}]$, the results support that improvements are expected as this term deviates from 1 (i.e., decreasing $\gamma$ or increasing $\Delta_{\mu}$).